# Structural Anisotropy Stabilises Asymmetric Beating in Instability Driven Filaments

### v1

@@ Appendix A Initial buckling for $\kappa_{0}\neq 0$
$\tau_{\perp}=-K_{B}\kappa_{0}||(\bm{I}-\hat{\bm{y}}\hat{\bm{y}})\hat{\bm{\mu}}||.$ (15)
In essence, the filament now has a now non-planar intrinsic curvature, causing off-plane torques which yield an earlier buckling. The effect of these torques is amplified as we increase $\kappa_{0},$ explaining why the filament buckles earlier as we increase the intrinsic curvature.
In Figure 9b, we consider these off-plane effects by calculating the in-plane contributions of $\hat{\bm{\mu}},$ and hence of the intrinsic curvature. This data is taken from initial value problems in the case where the intrinsic curvature is defined in the material frame, i.e. using Eq 13. The initial filament configuration is given by the steady state, as found by JFNK, with an additional small perturbation of $\sim O(10^{-3}).$ We let this simulation run until periodic oscillations are achieved. As predicted by the stability analysis, the perturbations grow faster as $f$ increases, leading to quicker saturation into the periodic beating we call P1. We can contrast this with the lab-frame intrinsic curvature, where the internal moment is defined by Eq 14, in which the off-plane contribution of the torque is always $\tau_{\perp}=-K_{B}\kappa_{0}||(\bm{I}-\hat{\bm{y}}\hat{\bm{y}})\hat{\bm{y}}||=0.$
@@ B.1 Isotropic filaments with intrinsic curvature
The bifurcation diagram for $\kappa_{0}=0$, i.e. the symmetric case, is shown in Figure 4a. As discussed in [37], we see that the filament buckles at $f=f^{*}$, matching the planar case. At this point, both planar beating and whirling branches are born through a double Hopf bifurcation. Planar beating is initially unstable, and whirling initially stable. At $f\approx 137.3$, whirling becomes unstable through a supercritical Hopf bifurcation in which QP1 is born. At $f\approx 140.0$, a pitchfork bifurcation occurs, in which the QP1 branch terminates and planar beating becomes stable.
@@ B.2 Anisotropic filaments with zero-curvature rest state
Appendix C Supplementary videos

### v2

@@ Appendix A Initial buckling for $\kappa_{0}\neq 0$
[…]
In essence, the filament now has a non-planar intrinsic curvature, causing off-plane torques which yield an earlier buckling. The effect of these torques is amplified as we increase $\kappa_{0},$ explaining why the filament buckles earlier as we increase the intrinsic curvature.
[…]
@@ B.1 Isotropic filaments with intrinsic curvature
The bifurcation diagram for $\kappa_{0}=0$, i.e. the symmetric case, is shown in Figure 4a. As discussed in (Clarke2024BifurcationsFilaments, ), we see that the filament buckles at $f=f^{*}$, matching the planar case. At this point, both planar beating and whirling branches are born through a double Hopf bifurcation. Planar beating is initially unstable, and whirling initially stable. At $f\approx 137.3$, whirling becomes unstable through a supercritical Hopf bifurcation in which QP1 is born. At $f\approx 140.0$, a pitchfork bifurcation occurs, in which the QP1 branch terminates and planar beating becomes stable.
@@ B.2 Anisotropic filaments with zero-curvature rest state
Figure 11: (a) The frequencies, $\omega=1/T$ where $T$ is the solution period obtained via JFNK, of three solution branches for $f\in[50,300]$. We show asymmetric beating for $(\kappa_{0},\beta)=(3\pi/4,1),$ P1 for $(\kappa_{0},\beta)=(3\pi/4,1),$ and whirling for $(\kappa_{0},\beta)=(\pi/4,1).$ The frequencies for various $\kappa_{0}$ and $\beta$ values for (b) whirling, (c) asymmetric beating and (d) P1 solutions are also shown for $f\in[200,210],$ to show the small variations in frequency as these parameters change.
Appendix C Solution Frequencies
In Figure 11a, we plot the frequencies, $\omega=1/T$, of some of the key dynamics observed in the main text; asymmetric beating, P1, and whirling. These are each plotted for a single choice of $\kappa_{0}$ and $\beta$ for $f\in[50,300]$. While the frequency of the solutions vary considerably with the follower force, as shown in Figure 11a, the change with the other parameters is much smaller, as shown in Figures 11b-d. The frequency of the whirling solutions increases with $\kappa_{0}$ and decreases with $\beta$ (Figure 11b) while the frequency of the asymmetric beating and P1 solutions both decrease with $\kappa_{0}$ (Figures 11c and d, respectively).
Appendix D Supplementary videos